OpenStudy (anonymous):

csc(x) = (4√3)/(3) and cos(x)<0. Evaluate the function sin(x). I know that sinθ = O/H and I know that H = 4√3 and A = 3. How do I find O?

jimthompson5910 (jim_thompson5910):

you don't need to do that since csc(x) = 1/sin(x)

jimthompson5910 (jim_thompson5910):

or sin(x) = 1/csc(x)

OpenStudy (campbell_st):

this is sin positive cos negative which is in the 2nd quadrant but the easy way to do the problem is to recognise cosec(x) = 1/sin(x) so sin(x) = 1/cosec(x) \[\sin(x) = \frac{3}{4\sqrt{3}}\]

OpenStudy (campbell_st):

so since the 2nd quadrant the angle will be 180 - x

OpenStudy (anonymous):

Thanks! That was really helpful!

OpenStudy (campbell_st):

good luck

OpenStudy (anonymous):

Wait! Small problem... that's not one of my possible answers. Give me a second and I'll post them.

jimthompson5910 (jim_thompson5910):

chances are you have to rationalize the denominator to get \[\Large \frac{3\sqrt{3}}{4}\]

OpenStudy (anonymous):

That's still not one :(

jimthompson5910 (jim_thompson5910):

alright post your possible answers please

jimthompson5910 (jim_thompson5910):

oh wait, I made a typo lol...It should be \[\Large \frac{3\sqrt{3}}{12}\]

jimthompson5910 (jim_thompson5910):

which reduces to \[\Large \frac{\sqrt{3}}{4}\]

OpenStudy (anonymous):

A. sin(x) = -(√13)/(4) B. sin(x) = -(4√13)/(13) C. sin(x) = (√3)/(4) D. sin(x) = -(√39)/(13) E. sin(x) = -(√39)/(3)

jimthompson5910 (jim_thompson5910):

So choice C

jimthompson5910 (jim_thompson5910):

Let me know if you need help rationalizing the denominator

OpenStudy (anonymous):

Thanks! I will :)

jimthompson5910 (jim_thompson5910):

Basically, to rationalize the denominator, you multiply both top and bottom of the fraction by sqrt(3). This will make \[\Large \frac{3}{4\sqrt{3}}\] turn into \[\Large \frac{3\sqrt{3}}{4\sqrt{3}\sqrt{3}}\] which becomes \[\Large \frac{3\sqrt{3}}{12}\] After fully simplifying, we get \[\Large \frac{\sqrt{3}}{4}\] So \[\Large \sin(x) = \frac{\sqrt{3}}{4}\]